OpenStudy (anonymous):

perform the indicated function evaluations for f(z)=squareroot 4z+5 a)f(0)

OpenStudy (anonymous):

For \(f(0)\) put z = 0 in the given..

OpenStudy (cwrw238):

plug in 0 for z in f(z)

OpenStudy (anonymous):

\[f(0) = \sqrt{4(0) + 5} = \sqrt{5}\]